OpenStudy (kkutie7):

This deals with t- test. d. For each method, is the Aluminum content different at 95% confidence level for the two Lots? Which technique offers the superior precision? I've done majority of the work for this and will provide certain answers if requested. I don't know how to use the t-table to answer this.

OpenStudy (kkutie7):

https://controls.engin.umich.edu/wiki/images/0/0a/DOF.jpg this is the t-table I'm supposed to use I guess

OpenStudy (perl):

what are parts a,b,c, , i need more information

OpenStudy (kkutie7):

it's just the answers for mean, standard deviation, spooled, and t-calc.

OpenStudy (kkutie7):

I wanna know how to use the table and the other things I've listed to answer the question.

OpenStudy (anonymous):

lololol okay was so lost. Just saw the link, give me a sec to think :)

OpenStudy (kkutie7):

alright

OpenStudy (kkutie7):

@kirbykirby @zepdrix

OpenStudy (anonymous):

I'm sorry.... I'm trying to figure out what you're asking, but you're using terms that I don't recognize. Your question, if it is a simple math problem, is worded in a very bizarre way.... :( I'm really sorry

OpenStudy (kkutie7):

It's analytical chemistry but can be counted at statistics which is math oriented.

OpenStudy (kirbykirby):

I think you should post the entire question... there is not enough being said from part d) only to know what the context of the problem is.

OpenStudy (kirbykirby):

You say "for each method" for example.. we don't even know what these methods are

OpenStudy (kkutie7):

alright. 4. An analyst was using two different methods (I and II) to monitor the content of Al in two different lots of an Aluminum nitride power (AIN) that his company was producing. Method I involved the use of a gravimetric method to determine the % Al in the AIN. The % Al results from Method II were obtained using an atomic emission spectroscopic method called inductively coupled plasma (ICP). Method I: Lot 1: 65.27% Lot 2: 65.98% 65.22% 65.80% 65.44% 65.67% 65.50% 65.75% Method II: Lot 1: 65.5% Lot 2: 66.0% 65.1% 65.5% 65.4% 66.5% 65.8% 65.7%

OpenStudy (kkutie7):

a. Calculate the mean and standard deviation for each lot in each of the two methods. b. Calculate the pooled standard deviations for two lots for each method (I and II). c. Calculate the student’s t values (talc) for each method (I and II).

OpenStudy (perl):

calculate the student's what ?

OpenStudy (perl):

hey you can upload the whole question, use a screenshot or

OpenStudy (perl):

you dont have to type it, just take a picture with a camera

OpenStudy (kirbykirby):

Do you know how to set up a 95% confidence interval? You just need to apply the formula. I'm guessing you are using the pooled standard deviation as well for this. If we just doing a hypothesis test that the 2 means are different, then you can use the confidence interval: \((\bar{x_1}-\bar{x_2})\pm t_{\text{crit}}*(S_p\sqrt{1/n_1+1/n_2}) \) , where S_p your pooled std. deviation and \(t_{crit}\) is the value you look for in your t-table. So look at the 95% column, and then the degrees of freedom corresponds to the sample size total - 2, so \(4+4-2\)

OpenStudy (kirbykirby):

there did you get 7.1 for your critical t value?

OpenStudy (kkutie7):

I used the t calc i calculated... I'm guess that was wrong it should have been the first number I wrote from the t-table?

OpenStudy (kirbykirby):

t=2.447

OpenStudy (kkutie7):

shoot you are right so this then? \[(65.63-65.80)\pm 2.447*(.00145\sqrt{1/4+1/4})\] \[-.17\pm.00355\sqrt{.5}\] \[-.17\pm.00251\]

OpenStudy (kirbykirby):

I believe so, assuming the other numbers are correct from a)...c)

OpenStudy (kirbykirby):

And the better method, I believe, should give a smaller confidence interval

OpenStudy (kirbykirby):

well the one offering more precision

OpenStudy (kkutie7):

ok but how do i see if they are different?

OpenStudy (anonymous):

Sorry to interject, but are we sure we're comparing means and not proportions? The data seems to indicate the latter, but I don't know for sure.

OpenStudy (kirbykirby):

Ohh I think you might be right on that one actually. Hm my brain is really tired tonight after that probability problem xD

OpenStudy (kkutie7):

I don't know =/ its been awhile since I've done work like this

OpenStudy (kkutie7):

answer to a: Method I Mean Standard Deviation Lot 1 65.63 .001158 Lot 2 65.80 .001138 Method II Mean Standard Deviation Lot 1 65.45 0.002887 Lot 2 65.93 0.004349

OpenStudy (kkutie7):

answer to b: Method I: \[pooled=√(((4-1) 〖.001158〗^2+(4-1) 〖.001138〗^2)/(4+4-2))\] \[=√((3(〖.001158〗^2 )+3(〖.001138〗^2))/6)\]\[=√((3(.00000134)+3(.00000130))/6)\]\[=√.00000132=.00145\] Method II: \[pooled=√(((4-1) 〖.002887〗^2+(4-1) 〖.004349〗^2)/(4+4-2))\] \[=√((3(〖.002887〗^2 )+3(〖.004349〗^2))/6)\]\[=√((3(.00000833)+3(.0000190))/6)\]\[=√.0000137=.00370\]

OpenStudy (kkutie7):

and c: Method I: \[t_calc=|65.80-65.63|/√(.001158/4+.001138/4)\] \[=|.17|/√(.0002895+.0002845)\] \[=.17/√.000574=.17/.023958=7.1\] Method II: \[t_calc=|65.93-65.45|/√(0.002887/4+0.004349/4)\] \[=|.48|/√(.000722+.001087)\] \[=.48/√.001809=.48/.042532=11.3\]

OpenStudy (kkutie7):

more information...

OpenStudy (perl):

whats up with the question marks?

OpenStudy (kkutie7):

noticed it tends to stay like that when I first input it but afterawhile it goes back to normal on from what I see. No question marks on my screen. @perl are you any good at chemistry?

OpenStudy (perl):

yes a bit

OpenStudy (perl):

whats the q.